solve for y: 24y +11 sqrt of y -18=0
Answer by pushpaharan(47)

let sqrt of y =x
24x^2+11x-18 =0
(8x+9)(3x-2)=0
x= -9/8 or x= 2/3
therfore y = 81/64 or y= 4/9